jester estimate that the probability that he will pass statistics and probability is 65%; the probability that he will pass basic calculus is 52%; and the probability of passing both subjects is 45%. find the probability that he will pass at least one of these two subjects



1
Expert's answer
2022-04-06T15:34:50-0400

Let A be the event of passing statistics and probability,

B - the event of passing basic calculus,

A "\cap" B - the event of passing both subjects.

The probability of pass at least one of these two subjects:

"P(A\cup B) =P(A) +P(B) - P(A\cap B) =\\=0.65+0.52-0.45=\\=0.72=72\%."